OpenStudy (anonymous):

An isosceles right triangle has it's area 20x².Find the area of the equilateral triangle constructed on one of it's legs ?

hartnn (hartnn):

so can you first find the measure of one of the leg ?

hartnn (hartnn):

A = 1/2 * leg^2 = 20 x^2 leg = ..?